1. Find the length of the curve x = 13 – 6t2, y = + + 6t2, t€ (0,1).
Answer #1

The length of u= f(t) , y = g(t) enterval [a, b] is given by length of a curve of the form the closed on t=6 7. 1)? It t=a WeThe answer sheet has three pages.it is the first pageceine on Thus, the length of the given [0, 1] is given by V (92 - 724 + 144+²)+ (a1 + 722 +1442%24 t=1 3. t=0 ✓ 18t² + 14422second page3 8 3 sont to 32 VTS [zd? [2] 10 (8) ntl r8 3√2 3 √2 [27 - 8x88 = V2 [27 - 8x252 272-32 6.1837 - 6.18 Korrect cepto two decimThird/last page